# Non-Overlap-Aware Egocentric Pose Estimation for Collaborative Perception in Connected Autonomy

### v1

@@ IV-A Experimental Setup
1) Precision is defined as the ratio of correctly retrieved correspondences to the retrieved correspondences, 2) Recall is defined as the ratio of correctly retrieved correspondences to the ground truth correspondences, 3) F1-Score, which evaluates the overall performance of the CoID method, is calculated as F1=$\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, 4) Position Error (PE) measures the Euclidean distance between the estimated and the ground truth position, 5) Rotation Error (RE) is defined as the geodesic distance [39] between the estimated rotation and the ground truth position, 6) Packet Size ((PS)) refers to the size of the data transmission packets shared between connected vehicles, to evaluate communication efficiency, 7) Non-Overlapping Detection Accuracy (NDA)) is defined as the ratio of correctly detected non-overlapping observation pairs to the total number of observation pairs, to evaluate accuracy of non-overlapping identification.

### v2

@@ IV-A Experimental Setup
1) Precision is defined as the ratio of correctly retrieved correspondences to the retrieved correspondences, 2) Recall is defined as the ratio of correctly retrieved correspondences to the ground truth correspondences, 3) F1-Score, which evaluates the overall performance of the CoID method, is calculated as F1=$\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, 4) Position Error (PE) measures the Euclidean distance between the estimated and the ground truth position, 5) Rotation Error (RE) is defined similarly to the rotation loss, where the quaternion and the ground truth quaternion is used as a surrogate for the geodesic distance [39], 6) Packet Size ((PS)) refers to the size of the data transmission packets shared between connected vehicles, to evaluate communication efficiency, 7) Non-Overlapping Detection Accuracy (NDA)) is defined as the ratio of correctly detected non-overlapping observation pairs to the total number of observation pairs, to evaluate accuracy of non-overlapping identification.